# Inertia onset in disordered porous media flow

### v2

@@ Abstract
We investigate the very onset of the inertial regime in pore-scale fluid flow in a three-dimensional, disordered porous media. We analyze how the flow structure changes with increasing Reynolds number. In particular, we compute tortuosity, spatial kinetic energy localization, and the pore-space volume fraction containing negative streamwise velocity. Our analysis shows that the onset of inertia observed in a standard way by computing the friction factor appears at a Reynolds number two orders of magnitude higher than indicated by analyzing tortuosity and spatial distribution of kinetic energy.
@@ I Introduction
One of the most striking phenomena in transport through porous media is the formation of flow channels, determined not only by geometry but by the transport phenomenon itself [1]. Several factors influence fluid behavior in porous media, such as pore size distribution, porosity, permeability, and the complexity of the fluid’s pathways, often referred to as tortuosity [2]. Its understanding is fundamental in hydrology, petroleum engineering, environmental science, and chemical engineering. Applications of porous media span across various fields, including microfluidic devices [3], energy production and development [4], and the design of filters [5] and chemical reactors [6], just to name a few. Usually, there are two empirical laws used to describe fluid dynamics in porous media. In creeping, laminar regime, [7, 8] Darcy’s law $\mathbf{u}=-\frac{\mathcal{K}}{\mu L}\nabla p$ is used, where $\mathbf{u}$ is the velocity vector, $\mu$ the dynamic viscosity of the fluid, $\mathcal{K}$ the permeability of the domain, $L$ the characteristic length of sample, and $p$ the pressure. However, the deviation from linear Darcy’s law emerges for higher velocities. There, the nonlinear relationship between the pressure drop and the filtration velocity is described by introducing the Forchheimer term, which is the second order correction to Darcy law: $\mathbf{u}=-\frac{\mathcal{K}}{\mu L}\nabla p-\beta\mathcal{K}|\mathbf{u}|\mathbf{u}$, [9, 10, 11, 12], where $\beta$ is a model parameter scaling the quadratic velocity term. The impact of large-scale permeability heterogeneity on high-velocity flows in porous media was analyzed in [13]. There, an effective inertial coefficient, which was determined from numerical simulations, varied with Reynolds number, variance, and mean permeability ratio differently across serial-layered, parallel-layered, and correlated media.
@@
The boundaries between Darcy and inertial flow regimes in two-dimensional porous media were experimentally observed for the first time in the rod porous medium model [14]. It was shown that the transition between linear and nonlinear regimes occurs for $101<Re<$ 10. In two-dimensional, disordered porous media, $Re=0.37$ was indicated [15]. A study utilizing network modeling to analyze the onset of non-Darcy flow in porous media highlighted the significant influence of pore geometry on the critical Reynolds number and the onset of inertial flow [16]. The influence of high fluid velocities on fluid flow and tracer transport in heterogeneous porous media was investigated in [17]. Structural complexity in the media was shown to intensify inertial effects and lead to a more homogenized flow [18]. In consequence, predictable chemical transport patterns were observed.
@@ III Results and Discussions
First, we identify the flow deviation between Darcy and the inertial regime by calculating the friction factor ($f\equiv-\Delta p/D\beta\rho u^{2}$) [15] for various Reynolds numbers, shown in Fig. 1. The results are presented as averages from eight independent samples along with the standard error of the mean. This result shows the regime where Darcy’s law correctly describes the system. Here, the transition regime appears between $1\leq Re\leq 10$.
@@
The results of $\varrho^{-}$ are illustrated in Fig. 3. Interestingly, the Reynolds number at which the backward flow emerges ($Re\approx 10$) is not compliant with Reynolds at which inertia appears, indicated in Fig. 1 and Fig. 2. To understand this, we calculate the spatial distribution of kinetic energy within the porous samples, a statistical measure of the channeling effect [15]. To quantify this phenomenon, we compute the energy localization defined as participation number $\pi$ [15]:
@@
Figure 4: Participation number ($\pi$) and its derivative ($\pi^{\prime}=d\pi/dlog_{10}Re$) (inset).
@@ IV Conclusions
In that study, we investigated what physical phenomena drive the transition from Darcy to the inertial regime in the flow through disordered porous media. Based on the results, we conclude that the physical inertial effects in the flow appear earlier than deviations from the Darcy law. Analysis of tortuosity $\tau$ and spatial energy distribution $\pi$ show visible impact of inertial forces on the flow for Re $\approx 1$. Moreover, derivatives of $\tau^{\prime}$ and $\pi^{\prime}$ are non-zero already at Re $\approx 0.1$. The weakening of the channeling effect with increasing $Re$ is also visible for $Re>1$ in streamline visualization. On the other hand, vortices appear at $Re\approx 10$ as indicated by the emergence of the backward flow in pore space. At approximately the same $Re$, the maximum of $\pi^{\prime}$ is observed. This indicates that vortices in pore space limit the kinetic energy dispersion in the system by excluding part of the pore space volume available for the main transport.

### v3

@@ Abstract
We investigate the very onset of the inertial regime in the fluid flow at the pore level in a three-dimensional, disordered, highly porous media. We analyze the flow structure in a wide range of Reynolds numbers starting from 0.01 up to 100. We focus on qualitative and quantitative changes that appear with increasing Reynolds number. To do that, we investigate the weakening of the channeling effect, defined as the existence of preferred flow paths in a system. We compute tortuosity, spatial kinetic energy localization, and the pore-space volume fraction containing negative streamwise velocity to assess accompanying changes quantitatively. Our results of tortuosity and participation number derivatives show that the very onset of inertia is apparent for Reynolds number Re $\sim 0.1$, an order of magnitude lower than indicated by analyzing relations of friction factor with the Reynolds number. Moreover, we show that the vortex structures appear at Reynolds number two orders of magnitude higher than the onset of inertia.
@@ I Introduction
One of the most striking phenomena in transport through porous media is the formation of flow channels, determined not only by the pore geometry but by the transport phenomenon itself [1]. Similar phenomena occurs in two-phase fluid flow as shown experimentally in [2, 3]. Several factors influence fluid behavior in porous media, such as pore size distribution, porosity, permeability, and the complexity of fluid pathways, often referred to as tortuosity [4]. Its understanding is fundamental in hydrology, petroleum engineering, environmental science, and chemical engineering. Applications of porous media span across various fields, including microfluidic devices [5], energy production and development [6], and the design of filters [7] and chemical reactors [8], just to name a few. Usually, there are two empirical laws used to describe fluid dynamics in porous media. In creeping, laminar regime, [9, 10] Darcy’s law $\mathbf{u}=-\frac{\mathcal{K}}{\mu L}\nabla p$ is used, where $\mathbf{u}$ is the velocity vector, $\mu$ the dynamic viscosity of the fluid, $\mathcal{K}$ the permeability of the domain, $L$ the streamwise length of porous sample, and $p$ the pressure. However, the deviation from linear Darcy’s law are observed for lower [11, 12] and higher velocities. In the latter, the nonlinear relationship between the pressure drop and the filtration velocity (or flowrate) is described by introducing the Forchheimer term, which is the second order correction to Darcy law: $\mathbf{u}=-\frac{\mathcal{K}}{\mu L}\nabla p-\beta\mathcal{K}|\mathbf{u}|\mathbf{u}$, [13, 14, 15, 16], where $\beta$ is a model parameter scaling the quadratic velocity term. The impact of large-scale permeability heterogeneity on high-velocity flows in porous media was analyzed in [17]. There, an effective inertial coefficient, which was determined from numerical simulations, varied with Reynolds number, variance, and mean permeability ratio differently across serial-layered, parallel-layered, and correlated media.
@@
The boundaries between Darcy and inertial flow regimes in two-dimensional porous media were experimentally observed for the first time in the rod porous medium model [18]. It was shown that the transition between linear and nonlinear regimes occurs for $101<Re<$ 10. Similar range of Reynolds Number $9.20.83<Re<$ 9.2 was reported in [19], where fluid flow through five different, three dimensional, sand samples was analyzed experimentally. Forchheimer’s flow through four different permeable stones’ samples was investigated and the influence of porosity, mean pore size and pore size distribution on transition to non-Darcy flow was discussed in  [20]. In two-dimensional, disordered porous media, $Re=0.37$ was indicated [21]. A study utilizing network modeling to analyze the onset of non-Darcy flow in porous media highlighted the significant influence of pore geometry on the critical Reynolds number and the onset of inertial flow [22]. The influence of high fluid velocities on fluid flow and tracer transport in heterogeneous porous media was investigated in [23]. An experimental study of two-phase flow through a porous sandstone sample reports the existence of three flow regimes. The authors detect the onset of inertia between the capillary-dominated and intermittent flow regimes [24]. Structural complexity in the media was shown to intensify inertial effects and lead to a more homogenized flow [25]. In consequence, predictable chemical transport patterns were observed.
@@ III Results and Discussions
First, we identify the flow deviation between Darcy and the inertial regime by calculating the friction factor ($f\equiv-\Delta p/D\beta\rho u^{2}$) [21] for various Reynolds numbers, shown in Fig. 1. The results are presented as averages from eight independent samples along with the standard error of the mean. This result shows the regime where Darcy’s law correctly describes the system. Here, the transition regime appears between $1\leq Re\leq 10$, which is in agreement with experimental results [19].
@@
The results of $\varrho^{-}$ are illustrated in Fig. 3. Interestingly, the Reynolds number at which the backward flow emerges ($Re\approx 10$) is not compliant with Reynolds number at which inertia appears, indicated in Fig. 1 and Fig. 2. To understand this, we calculate the spatial distribution of kinetic energy within the porous samples, a statistical measure of the channeling effect [21]. To quantify this phenomenon, we compute the energy localization defined as participation number $\pi$ [21]:
@@
Figure 4: Participation number ($\pi$) and its derivative ($\pi^{\prime}=d\pi/d\mathrm{log}_{10}Re$) (inset).
@@ IV Conclusions
In that study, we investigated what physical phenomena drive the transition from Darcy to the inertial regime in the flow through disordered porous media.
By examining Reynolds numbers from $0.01$ to $100$, we identify significant qualitative and quantitative changes in the flow structure as inertia becomes significant. Our findings indicate that the onset of inertia, characterized by a weakening of the channeling effect and the appearance of vortices, occurs at a Reynolds number ($Re$) as low as 0.1. This is an order of magnitude lower than the traditionally accepted value derived from the friction factor-Reynolds number relationship. This early onset of inertial effects is marked by changes in tortuosity and the spatial distribution of kinetic energy, with tortuosity beginning to decrease at $Re\approx 1$. The study also reveals that the volume fraction of the pore space containing negative streamwise velocity emerges at $Re\approx 10$, indicating the formation of backward flows and vortices, which become more pronounced at higher Reynolds numbers. These phenomena are further confirmed by the increase in the participation number, suggesting a more homogeneous energy distribution and a reduction in the channeling effect. The visualization of streamlines supports this conclusion, showing a transition from highly channeled to more dispersed flow patterns as $Re$ increases. The emergence of backward flow and the resulting vortices at higher $Re$ indicate complex flow interactions, hinting at the precursor to turbulence in porous media. Our study highlights the nuanced and early onset of inertial effects in disordered porous media, providing a deeper understanding of the transition from Darcy to non-Darcy flow regimes. Our findings underscore the importance of considering pore-scale flow dynamics in studying fluid flow in the inertial regime in porous systems.